# Decentralized Nonconvex Optimization under Heavy-Tailed Noise: Normalization and Optimal Convergence

### v1

@@ 4 Main Results
We present the main convergence results of GT-NSGDm  and discuss their implications. The detailed analyses are deferred to the Appendix B. We first consider the case where the tail index $p$ is known.
@@
Let Assumptions 1, 2, 3, 4 hold. Denote $f(\bar{\bm{x}}^{0})-f_{*}=\Delta_{0},[\nabla f_{1}(\bar{\bm{x}}^{0}),\ldots,\nabla f_{n}(\bar{\bm{x}}^{0})]^{\top}=\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})$. Take
$\alpha=\min\Big{(}1,\ \max\Big{(}\sqrt{\frac{\Delta_{0}(1-\beta)(1-\lambda)}{4LT}},\sqrt{\frac{\Delta_{0}(1-\lambda)}{(1+2.5L)T}},\sqrt{\frac{(1-\lambda)^{2}\Delta_{0}}{2n^{\frac{1}{2}}LT}}\Big{)}\Big{)},$ (5)
and $1-\beta=1/T^{\frac{p}{3p-2}}$. Assume $\beta\geq 1/10$, then the sequence generated from GT-NSGDm satisfies that
$\frac{1}{nT}\sum_{t=0}^{T-1}\sum_{i=1}^{n}\mathbb{E}\big{[}\|\nabla f(\bm{x}_{i}^{t})\|]=O\Big{(}\frac{\sigma}{n^{1-\frac{1}{p}}T^{\frac{p-1}{3p-2}}}+\frac{1}{T^{\frac{p-1}{3p-2}}}\sqrt{\frac{L\Delta_{0}}{1-\lambda}}+\frac{\|\nabla f(\bar{\bm{x}}^{0})\|}{T^{\frac{2p-2}{3p-2}}}+\sqrt{\frac{(1+2.5L)\Delta_{0}}{(1-\lambda)T}}$
@@
We discuss the asymptotic speedup in number of nodes $n$. For sufficiently large $T$ (or sufficiently small target optimality gap), the upper bound in Theorem 1 is dominated by the leading terms
$(1/T^{\frac{p-1}{3p-2}})\Big{(}\frac{\sigma}{n^{1-1/p}}+\sqrt{\frac{L\Delta_{0}}{1-\lambda}}\Big{)}.$
In the high-noise regime $\sigma\gg n^{1-\frac{1}{p}}\sqrt{\frac{L\Delta_{0}}{1-\lambda}}$, the upper bound has a speedup factor $n^{1-\frac{1}{p}}$. In practice, the noise scale (measured by $\sigma$) in training attention models or in other high-dimensional problems can be very large, and the speedup in $n$ contributes as a noise reduction.
@@
[…]
$\frac{\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{(1-\lambda)n^{\frac{1}{2}}}\cdot\frac{1}{\sqrt{T}}+\sqrt{\frac{n^{\frac{1}{2}}L\Delta_{0}}{(1-\lambda)^{2}T}}+\frac{\sigma n^{\frac{1}{2}}}{1-\lambda}\cdot\frac{1}{T^{\frac{2p-1}{2p}}}+\frac{\Delta_{0}}{T}\Big{)}.$
@@ 5 Experiments
We assess the performance of GT-NSGDm through numerical experiments. We first conduct empirical studies on synthetic datasets that mimic language modeling under controlled heavy-tailed noise injection, following a similar setup in [35]. We also present experimental studies on decentralized training of a small decoder-only Transformer (GPT) model with 3M parameters on the Multi30k dataset.
Baseline methods. We compare GT-NSGDm with several decentralized baseline methods: DSGD [55], DSGD-Clip [29], GT-DSGD [7], and SClip-EF-Network [30]. Among these, DSGD and GT-DSGD serve as baselines for decentralized optimization under regular stochastic noise with bounded variance. DSGD-Clip with decaying stepsize and increasing $\ell_{2}$ clipping levels, is shown to converge for strongly convex functions assuming compact domain in [56]. SClip-EF-Network achieves convergence under symmetric noise with bounded $\mathbb{E}\big{[}\|\bm{\xi}_{i}^{t}\|^{p}\mid\mathcal{F}_{t-1}\big{]}$ for $p=1$. All baseline methods are initialized from the same point. A detailed summary of these baselines is provided in Table 1 in Appendix C.1. In our experiments, all algorithms are tuned from grid search.
Network Topology. We consider the three different graph topologies: undirected ring graphs, directed exponential graphs, and complete graphs (See [6, 2, 54] for details on graph configurations). For each topology, we compute the mixing matrix using Metropolis weights [57]. In our synthetic experiments, we set the number of nodes to $n=20$, resulting in $\lambda=0.904$, $0.714$, and 00 for the ring, exponential, and complete graphs, respectively. For decentralized training of Transformer models, we use $n=8$, with corresponding $\lambda=0.804$, $0.6$, and 00.
@@ 5.1 Robust linear regression on synthetic tokenized data
We consider a nonconvex regularized linear regression model, on a synthetic dataset that mimics language tokens. In language modeling, token frequencies often exhibit a heavy-tailed distribution: a few tokens appear very frequently, while the vast majority are rare but contribute essential contextual meaning. We construct the following synthetic dataset $\bm{X}$ of 1k samples of dimension $d=20$. The first two features are to simulate frequent tokens, and are sampled from Bernoulli distributions Bern(0.9) and Bern(0.5), respectively. The remaining 8 features represent rare tokens and are each sample from Bern(0.1). The optimal weight vector $\bm{w}_{*}$ is sampled from a Gaussian distribution, and true label vector is $\bm{y}=\bm{X}\bm{w}_{*}$. The synthetic dataset $(\bm{X},\bm{y})$ is evenly distributed over $n=20$ nodes, where each node $i$ holds a sub-dataset $(\bm{X}_{i},\bm{y}_{i})$, estimate $\bm{w}_{i}$, and a linear regression model with nonconvex robust Tukey’s biweight loss function [58] to estimate the ground truth parameter $\bm{w}_{*}$. We inject three different zero-mean noises, Gaussian noise ($\mathcal{N}(\mathbf{0},3\bm{I}_{d})$), Student’s $t$ noise (degrees of freedom $1.5$, scale $1.0$), and Lévy $\alpha$-stable noise (stability parameter $1.5$, skewness parameter $0.5$, scale $1.0$, non-symmetric, multiplied by $0.1$) into computed exact gradient, and use corrupted stochastic gradients for algorithm updates. See Appendix C.2 for additional details.
In Figure 2, we evaluate the performance of GT-NSGDm against baseline methods on a ring graph under various types of injected gradient noise. The results show that while DSGD and GT-DSGD converge under Gaussian noise (which has bounded variance), they become unstable when exposed to heavy-tailed noise. Although DSGD-Clip remains stable across all noise types, it fails to converge to the optimum. In contrast, both GT-NSGDm and SClip-EF-Network exhibit robust convergence behavior and achieve near-optimal performance consistently across all noise scenarios. Although SClip-EF-Network outperforms GT-NSGDm by a small margin in this low-dimensional synthetic problem, GT-NSGDm outperforms it significantly in Transformers training over a real-world corpus.
@@
In Figure 3, we study GT-NSGDm’s empirical dependence on problem parameters including connectivity ($\lambda$), noise level ($\sigma$), and the number of nodes ($n$). We vary each parameter while keeping the others fixed. In Figure 3(a), we inject the aforementioned Lévy $\alpha$-stable noise and test the performance of GT-NSGDm on ring, directed exponential, undirected exponential, and complete graphs with $\lambda=0.904,0.714,0.6,0$, respectively. It is shown that GT-NSGDm achieves comparable final errors under weak connectivity (i.e., large $\lambda$) comparing to complete graphs, demonstrating its favorable dependence on network connectivity under heavy-tailed noise. In Figure 3(b), we evaluate GT-NSGDm’s performance under different noise levels on a directed exponential graph. Under Gaussian noise with scale 1 (unit variance), GT-NSGDm reaches the best optimality; the final error increases as $\sigma$ grows, as observed under both Gaussian and Lévy $\alpha$-stable noise. In Figure 3(c), we inject Lévy $\alpha$-stable noise and use complete graphs ($\lambda=0$ for all $n$) with varying number of nodes. we observe that as $n$ increases from 2 to 40, convergence speed improves. The corresponding final errors, i.e., $[0.4,0.35,0.29,0.20,0.21]$, demonstrate a clear speedup trend over a certain range of $n$, supporting the theoretical speedup discussions in Remarks 5 and 6.
@@ 5.2 Decentralized training of Transformers
We further evaluate GT-NSGDm on real-world language modeling tasks. Specifically, we consider a GPT model [50] with approximately 3 million parameters and perform an auto-regressive language modeling task on the Multi30k dataset (29,000 sentences and 4.4 million tokens in the training split). We assess algorithm performance using validation logarithmic perplexity. On a graph of 8 nodes (with the three topologies), we evenly distribute the training corpus across the nodes and initialize an identical GPT model on each. We run GT-NSGDm and baseline methods for 12 epochs with a batch size of 64. See Appendix C.3 for additional details on model and hyperparameter search.
In Figure 4, we plot the average validation loss and its standard deviation over five independent runs for each algorithm across three different topologies. The results show that, in this high-dimensional real-world task, GT-NSGDm is the only algorithm that consistently converges to near-zero loss with low variance, achieving average final loss levels of 0.258, 0.261, and 0.282 on ring, directed exponential, and complete graphs, respectively. Other baseline methods, while stable, plateau at loss levels between 5 and 6. This sharp contrast underscores that GT-NSGDm is not only theoretically sound but also robust in real-world tasks.
(a) Ring Graph
(b) Directed Exponential Graph
(c) Complete Graph
Figure 4: Comparison of graph-wide average validation losses in decentralized training of Transformer models, over ring, directed exponential, and complete graphs.

### v2

@@ 4 Main Results
We present the main convergence results of GT-NSGDm and discuss their implications. The detailed analyses are deferred to Appendix B. We first consider the case where the tail index $p$ is known.
@@
Let Assumptions 1, 2, 3, 4 hold. Denote $f(\bar{\bm{x}}^{0})-f_{*}=\Delta_{0},[\nabla f_{1}(\bar{\bm{x}}^{0}),\ldots,\nabla f_{n}(\bar{\bm{x}}^{0})]^{\top}\\ =\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})$. Take
$\alpha=\min\Big(1,\ \sqrt{\frac{\Delta_{0}(1-\beta)(1-\lambda)}{4LT}},\sqrt{\frac{\Delta_{0}(1-\lambda)}{3.5LT}},\sqrt{\frac{(1-\lambda)^{2}\Delta_{0}}{2n^{\frac{1}{2}}LT}}\Big),$ (5)
and $1-\beta=1/T^{\frac{p}{3p-2}}$. Assume $\beta\geq 1/10$, then the sequence generated by GT-NSGDm satisfies that
$\frac{1}{nT}\sum_{t=0}^{T-1}\sum_{i=1}^{n}\mathbb{E}\big[\|\nabla f(\bm{x}_{i}^{t})\|]=O\Big(\frac{\sigma}{n^{1-\frac{1}{p}}T^{\frac{p-1}{3p-2}}}+\frac{1}{T^{\frac{p-1}{3p-2}}}\sqrt{\frac{L\Delta_{0}}{1-\lambda}}+\frac{\|\nabla f(\bar{\bm{x}}^{0})\|}{T^{\frac{2p-2}{3p-2}}}+\sqrt{\frac{3.5L\Delta_{0}}{(1-\lambda)T}}$
@@
We discuss the asymptotic speedup in number of nodes $n$. For sufficiently large $T$ (or sufficiently small target optimality gap), the upper bound in Theorem 1 is dominated by the leading terms $(1/T^{\frac{p-1}{3p-2}})\big(\sigma/n^{1-1/p}+\sqrt{L\Delta_{0}/(1-\lambda)}\big)$. In the high-noise regime $\sigma\gg n^{1-1/p}\sqrt{L\Delta_{0}/(1-\lambda)}$, the upper bound has a speedup factor $n^{1-1/p}$. In practice, the noise scale (measured by $\sigma$) in training attention models or in other high-dimensional problems can be very large, and the speedup in $n$ contributes as a noise reduction.
@@
$\frac{1}{nT}\sum_{t=0}^{T-1}\sum_{i=1}^{n}\mathbb{E}\big[\|\nabla f(\bm{x}_{i}^{t})\|]\leq O\Big(\frac{\sigma}{n^{1-\frac{1}{p}}T^{\frac{p-1}{2p}}}+\frac{1}{T^{\frac{1}{4}}}\sqrt{\frac{L\Delta_{0}}{1-\lambda}}+\frac{\|\nabla f(\bar{\bm{x}}^{0})\|}{\sqrt{T}}+\sqrt{\frac{3.5L\Delta_{0}}{(1-\lambda)T}}$
$+\frac{\sigma n^{\frac{1}{2}}}{(1-\lambda)^{\frac{1}{p}}\sqrt{T}}+\frac{\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{(1-\lambda)n^{\frac{1}{2}}\sqrt{T}}+\sqrt{\frac{n^{\frac{1}{2}}L\Delta_{0}}{(1-\lambda)^{2}T}}+\frac{\sigma n^{\frac{1}{2}}}{(1-\lambda)T^{\frac{2p-1}{2p}}}+\frac{\Delta_{0}}{T}\Big).$
@@
Remark 7 (Hyperparameter selection).
We compare with the centralized work [27] in terms of hyperparameter selections. When $p$ is known, the centralized setup requires $\Delta_{0},L,\sigma,T$, while our decentralized setup requires $\Delta_{0},L,T,\lambda,n$. Notably, our selection is independent of $\sigma$ but requires the network parameters $n,\lambda$. When $p$ is unknown, the centralized setup requires $L,T$, while the decentralized case additionally requires $\Delta_{0},\lambda,n$. We note that $n$ is generally easy to estimate, and we give an example where our hyperparameter dependence on $\lambda$ can be eliminated. If all nodes know $n$ and set weights $\bm{W}=\bm{I}_{n}-(1/n)\bm{L}$ where $\bm{L}$ is the graph Laplacian, one can obtain the estimate $\lambda\leq 1-c/n^{3}$ [55] for some universal constant $c$. Once we have such an upper bound, we can replace the terms involving $1/(1-\lambda)$ with this estimate in equation (35); then in the final proof steps, we can select a step size that is independent of $\lambda$ while still obtaining the same order-optimal rate.
Remark 8 (Dominant terms for small $T$).
Both Theorems 1 and 2 order the upper bound terms in an increasing order of the rate exponent of $1/T$, so the dominant terms are clear for sufficiently large $T$. We briefly discuss some cases to illustrate the effects of other parameters including $n$, $\lambda$, and $p$ when $T$ is small. We discuss one varying parameter while assuming other parameters are fixed. First, for network connectivity $1-\lambda$, $1/{T^{\frac{p-1}{3p-2}}}$ has coefficient $\frac{1}{\sqrt{1-\lambda}}$ while the faster decaying term $\frac{1}{\sqrt{T}}$ has coefficient $\frac{1}{1-\lambda}$. When $1-\lambda\leq\frac{n^{1/2}}{T^{p/(3p-2)}}$, the latter term $\frac{1}{\sqrt{T}}\sqrt{\frac{n^{1/2}L\Delta_{0}}{(1-\lambda)^{2}}}$ will dominate. Second, for $n$, if $n^{1/2}\geq(1-\lambda)T^{\frac{p}{3p-2}}$, we will also have a dominant term $\frac{1}{\sqrt{T}}\sqrt{\frac{n^{1/2}L\Delta_{0}}{(1-\lambda)^{2}}}$. Third, for $p\in(1,2]$: When $p$ approaches $1$ from above, $\frac{1}{T^{(p-1)/(3p-2)}}$ will have coefficient $\sqrt{\frac{L\Delta_{0}}{1-\lambda}}+\sigma$ and the speedup in terms of the number of nodes will vanish, but the change in $\frac{1}{T^{p/(3p-2)}}\frac{\sigma n^{1/2}}{(1-\lambda)^{1/p}}$ will not introduce a new dominant term if we focus on the dependence on $1-\lambda$ since there already exists a slower term $\frac{1}{\sqrt{T}}\frac{n^{1/2}}{1-\lambda}$. Similarly, when $p$ approaches $2$, the speedup in $n$ is more significant; the term $\frac{1}{T^{p/(3p-2)}}\frac{\sigma n^{1/2}}{(1-\lambda)^{1/p}}$ only gets smaller and does not change the dominant terms. Although our rates are established for general $p\in(1,2]$, we can compare parameter dependence with rates obtained for $p=2$ only. For example, in [56], for large enough $T$, the upper bound is
$O\Big(\frac{\Delta_{0}+\sigma\sqrt{L}}{n^{1/4}T^{1/4}}+\frac{\sqrt{n}\sigma L}{(1-\lambda^{2})^{3/2}\sqrt{T}}+\frac{L\|\nabla F(\mathbf{1}_{n}\otimes\bar{\bm{x}}^{0})\|}{(1-\lambda^{2})^{3/2}T}\Big).$
If we set $p=2$ in our bounds, dropping some higher order terms, we have
$O\Big(\frac{1}{T^{1/4}}\cdot\big(\sqrt{\frac{L\Delta_{0}}{1-\lambda}}+\frac{\sigma}{n^{1/2}}\big)+\frac{1}{\sqrt{T}}\cdot\big(\|\nabla f({\bar{\bm{x}}}^{0})\|+\sqrt{\frac{3.5L\Delta_{0}}{1-\lambda}}+\sqrt{\frac{n^{\frac{1}{2}}L\Delta_{0}}{(1-\lambda)^{2}}}+\frac{\sigma n^{\frac{1}{2}}}{(1-\lambda)^{1/2}}\big)\Big).$
We compare the above two bounds. For coefficients of $1/T^{1/4}$, our bound has dependence $1/n^{1/2}$ which is faster than the $1/n^{1/4}$ in the $p=2$ case, but the dependence in $1/\sqrt{1-\lambda}$ is worse. For coefficients of $1/\sqrt{T}$, our bounds have dependence $1/(1-\lambda)$ while the $p=2$ case has dependence $1/(1-\lambda^{2})^{3/2}$, which is strictly larger than ours.
@@ 5 Experiments
We assess the performance of GT-NSGDm through numerical experiments. We first conduct studies on synthetic datasets that mimic language modeling under controlled heavy-tailed noise injection, following [35]. We also present experiments on decentralized training of a decoder-only Transformer (GPT) model with 3M parameters on the Multi30k dataset.
Baselines. We compare GT-NSGDm with four decentralized baselines: DSGD[57], GT-DSGD [7], DSGD-Clip [29], and SClip-EF -Network [30]. DSGD and GT-DSGD handle regular stochastic noise with bounded variance. DSGD-Clip converges for strongly convex functions under bounded domains or gradients [29]. SClip-EF-Network achieves convergence under symmetric noise with bounded $\mathbb{E}\big[\|\bm{\xi}_{i}^{t}\|^{p}\mid\mathcal{F}_{t-1}\big]$ for $p=1$. All methods are initialized identically and tuned via grid search. Detailed baseline descriptions appear in Table 2 (Appendix C.1).
Graph topology. We consider three graph topologies: undirected ring, directed exponential, and complete graphs (see [6, 2, 54]). Weight matrices use Metropolis weights [58]. For synthetic experiments, we set the number of nodes to $n=20$, we obtain $\lambda=0.904$, $0.714$, and 0 for the ring, exponential, and complete graphs, respectively. For Transformer training with $n=8$, we have corresponding $\lambda=0.804$, $0.6$, and 0.
@@ 5.1 Robust linear regression on synthetic tokenized data
We use this synthetic experiment to test our convergence rates under controlled heavy-tailed noise. We consider nonconvex regularized linear regression on synthetic data mimicking language tokens. In language modeling, token frequencies exhibit heavy-tailed distributions: few tokens appear frequently, while most are rare but contextually important. We construct the following synthetic dataset $\bm{X}$ of 1k samples of dimension $d=20$. The first four features simulate frequent tokens and are sampled independently from Bernoulli distributions: the first two from $\mathrm{Bern}(0.9)$ and the next two from $\mathrm{Bern}(0.5)$. The remaining 16 features represent rare tokens, each sampled from $\mathrm{Bern}(0.1)$. The optimal weight $\bm{w}_{*}$ is Gaussian-sampled, with labels $\bm{y}=\bm{X}\bm{w}_{*}$. The synthetic dataset $(\bm{X},\bm{y})$ is evenly distributed over $n=20$ nodes, where each node $i$ holds a sub-dataset $(\bm{X}_{i},\bm{y}_{i})$, estimate $\bm{w}_{i}$, and a linear regression model with nonconvex robust Tukey’s biweight loss function [59] to estimate $\bm{w}_{*}$. We inject three different zero-mean noises, Gaussian noise ($\mathcal{N}(\mathbf{0},3\bm{I}_{d})$), Student’s $t$ noise (degrees of freedom $1.5$, scale $1.0$), and Lévy $\alpha$-stable noise (stability parameter $1.5$, skewness parameter $0.5$, scale $1.0$, non-symmetric, multiplied by $0.1$) into the exact gradient, using corrupted stochastic gradients for updates. See Appendix C.2 for additional details.
In Figure 2, we evaluate GT-NSGDm against baselines on ring graphs under various gradient noise. DSGD and GT-DSGD converge under Gaussian noise but become unstable under heavy-tailed noise. DSGD-Clip remains stable but fails to reach the optimum. Both GT-NSGDm and SClip-EF-Network exhibit robust convergence and near-optimal performance across all scenarios, consistent with their theoretical guarantees under heavy-tailed noise.
@@
In Figure 3, we test GT-NSGDm’s dependence on connectivity ($\lambda$), noise level ($\sigma$), and the number of nodes ($n$), varying each while fixing others. In Figure 3(a), we inject Lévy $\alpha$-stable noise and test the performance of GT-NSGDm on ring, directed exponential, undirected exponential, and complete graphs with $\lambda=0.904,0.714,0.6,0$, respectively. GT-NSGDm achieves comparable final errors under weak connectivity (i.e., large $\lambda$) versus complete graphs, showing favorable dependence on network connectivity under heavy-tailed noise. In Figure 3(b), we evaluate GT-NSGDm’s performance under different noise levels on a directed exponential graph. Under Gaussian noise with scale 1 (unit variance), GT-NSGDm reaches the best optimality; the final error increases as $\sigma$ grows, as observed under both Gaussian and Lévy $\alpha$-stable noise. In Figure 3(c), we inject Lévy $\alpha$-stable noise on complete graphs ($\lambda=0$ for all $n$) with varying number of nodes. As $n$ increases from 2 to 40, convergence speed improves with final errors $[0.4,0.35,0.29,0.20,0.21]$, demonstrating speedup over certain $n$ ranges, supporting theoretical discussions in Remarks 5 and 6.
@@ 5.2 Decentralized training of Transformers
We evaluate GT-NSGDm’s empirical performance on language modeling using a 3M-parameter GPT model [50] for auto-regressive modeling on Multi30k (29k sentences, 4.4M tokens). We assess performance using validation log-perplexity. On 8-node graphs with three topologies, we distribute training data evenly and initialize identical GPT models per node. We introduce three additional baselines: DSGD-GClip (DSGD with constant step size and $\ell_{2}$ gradient clipping level), DSGD-CClip (DSGD with constant step size and component-wise gradient clipping level), QG-DSGDm [60] and GT-Adam [61] (all without theoretical guarantees under heavy-tailed noise; see Table 2 in Appendix C.1). We run all methods for 12 epochs with batch size 64. See Appendix C.3 for model and hyperparameter details.
Table 1 presents average validation loss and standard deviation over five independent runs for each algorithm across three topologies. Results show that GT-NSGDm nearly matches the best baseline DSGD-GClip (which lacks theoretical guarantees under heavy-tailed gradient noise) while significantly outperforming the other two theoretically-guaranteed baselines across all topologies. We note that this decentralized training experiment is simulated to demonstrate algorithm effectiveness and has practical limitations.
Table 1: Topologies ring, directed exponential (Exp.), and complete (Comp.) graphs. Algorithms are grouped by theoretical (Theo.) guarantees under heavy-tailed noise: with (w/) or without (w/o).
Algorithms Theo. Ring Exp. Comp.
DSGD w/o $5.633_{\pm 0.008}$ $5.632_{\pm 0.007}$ $5.635_{\pm 0.007}$
DSGD-GClip w/o $\mathbf{0.253}_{\pm 0.007}$ $\mathbf{0.249}_{\pm 0.010}$ $\mathbf{0.267}_{\pm 0.010}$
DSGD-CClip w/o $2.725_{\pm 3.179}$ $5.058_{\pm 2.388}$ $8.225_{\pm 1.695}$
GT-DSGD w/o $5.362_{\pm 0.002}$ $5.632_{\pm 0.002}$ $5.631_{\pm 0.002}$
GT-Adam w/o $0.520_{\pm 0.038}$ $0.587_{\pm 0.096}$ $0.524_{\pm 0.045}$
QG-DSGDm w/o $0.394_{\pm 0.007}$ $0.388_{\pm 0.013}$ $0.353_{\pm 0.011}$
SClip-EF-Network w/ $5.653_{\pm 0.012}$ $5.632_{\pm 0.003}$ $5.636_{\pm 0.004}$
DSGD-Clip w/ $5.633_{\pm 0.004}$ $5.659_{\pm 0.013}$ $5.661_{\pm 0.006}$
GT-NSGDm w/ $\mathbf{0.258}_{\pm 0.007}$ $\mathbf{0.261}_{\pm 0.007}$ $\mathbf{0.282}_{\pm 0.009}$